OpenStudy (anonymous):

pls help with this, if r=d/2 find the area of the bigger semi-circle, subtract the area of the smaller semi-circle from it to get the area of the shaded region. 5in 10 in-shaded area

OpenStudy (anonymous):

so is this a 10-in semi-circle with a 5-in semi-circle inside it, and the area between the 5 and 10 are shaded?

OpenStudy (anonymous):

and the 10in and 5in are diameters?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it's the area of the bigger minus the area of the smaller. area of a semi circle is 1/2 pi*r^2. 1/2 pi*10^2 - 1/2 pi*2.5^2

OpenStudy (anonymous):

i meant 5 for the first radius, not 10. oops.

OpenStudy (anonymous):

but is that same as r=d/2

OpenStudy (anonymous):

ok im confused now could u go step by step listing

OpenStudy (anonymous):

OK, area of the larger semi circle is:\[A = \pi r^2\]and you know the diameter is 10, so the radius is 5. so:\[A=\pi 5^2\]similarly, the other circle is \[A = \pi 2.5^2\]The area of the shaded region is just the area of the larger minus the area of the smaller, so just calculate and subtract one from the other.

OpenStudy (anonymous):

ok so wat is my teacher tlk about using r=d/2 same thing rite?

OpenStudy (anonymous):

r=d/2 means radius equals half of diameter. it's a property of all circles.

OpenStudy (anonymous):

so okay with the steps given i shld find the answer

OpenStudy (anonymous):

yes, just multiply out\[\pi 5^2 - \pi 2.5^2 = 25\pi - 6.25\pi = 18.75\pi = 58.90\]

OpenStudy (anonymous):

thanku ok

OpenStudy (anonymous):

wait, i was calculating for full circles. the answer would be half for semicircles (29.45)

OpenStudy (anonymous):

is what i got a = pi(r^2) a = pi(5/2)^2 a = 25/4pi area larger circle is a = pi(r^2) a = pi(10/2)^2 a = pi(5^2) a = 25pi larger - smaller = 25pi - 25/4pi = 100/4pi - 25pi = 75pi/4